# Optimal Consumption and Portfolio Choice with No-Borrowing Constraint in the Kim-Omberg Model

### 03 Mar 2026 v1

@@ 4. Back to the Primal Problem
The process $\tau^{D}:=\{\tau^{D}(\alpha)\mid\alpha\in(0,1]\}$ has nonincreasing, left-continuous sample paths and hence it admits right-limits $\tau^{D}_{+}(\alpha):=\inf\{t\geq 0\mid D_{t}\leq\alpha\}$. The set of points $\alpha\in(0,1]$ at which $\tau^{D}(\alpha)(\omega)\neq\tau^{D}_{+}(\alpha)(\omega)$ is countable for a.e. $\omega\in\Omega$. Since $(D_{t})_{t}$ is right-continuous and $\tau^{D}(\alpha)$ is the first entry time of an open set, it is an $\bar{{\mathcal{F}}}_{t+}$-stopping time for any given and fixed $\alpha\in(0,1]$. However, $(\bar{\mathcal{F}}_{t})_{t}$ is right-continuous, hence $\tau^{D}(\alpha)$ is an $\bar{\mathbb{F}}$-stopping time. For a fixed $z>0$ and $y\in(0,z]$, we consider $\alpha=\frac{y}{z}$ in the following. Upon using sub-optimality of $\tau^{D}(\frac{y}{z})$ for (3.13), we have
@@
where we used that $D^{*}\in\mathcal{D}$ since $(\xi_{t}^{*})_{t}$ is nondecreasing, càdlàg, and $\bar{\mathbb{F}}$-adapted.
@@
$c_{t}^{*}=(Z_{t}^{\hat{z},\beta,D^{*}})^{-\frac{1}{\gamma}}\qquad\text{and}\qquad\pi_{t}^{*}=\frac{\beta_{t}}{\sigma^{2}}Z_{t}^{\hat{z},\beta,D^{*}}\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})-\frac{\rho\sigma_{\beta}}{\sigma}\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
@@
$=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(u(c_{t}^{*})-\tilde{u}(Z_{t}^{D^{*}})\big)\,dt\bigg]-\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\ell Z_{t}^{D^{*}}\,dt\bigg],$
and hence, by using (4.14), we obtain
@@
so that $c_{t}^{*}=(Z_{t}^{\hat{z},\beta,D^{*}})^{-\frac{1}{\gamma}}$ is optimal. By Proposition 2.6, there exists an investment strategy $\pi^{*}$ such that $(\pi^{*},c^{*})\in\mathcal{A}(x)$. Moreover, by the strong Markov property and $x=-\tilde{V}_{z}(\hat{z},\beta)$, we have
@@
$-\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\left[\kappa(\overline{\beta}-\beta_{t})\,dt+\sigma_{\beta}\,dW_{t}^{\beta}\right]$
@@
$-\frac{1}{2}\sigma_{\beta}^{2}\tilde{V}_{z\beta\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\,dt+\frac{\beta_{t}}{\sigma}Z_{t}^{\hat{z},\beta,D^{*}}\sigma_{\beta}\rho\tilde{V}_{zz\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\,dt,\quad t\geq 0.$
Recalling that $dW^{\beta}_{t}=\rho dW_{t}+\sqrt{1-\rho^{2}}dW^{2,\perp}_{t}$ due to $Corr(W_{t},W^{\beta}_{t})=\rho t$, comparing the $dW_{t}$ terms in (2.1) and (4.19) finally gives
$\pi_{t}^{*}=\frac{\beta_{t}}{\sigma^{2}}Z_{t}^{\hat{z},\beta,D^{*}}\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})-\frac{\rho\sigma_{\beta}}{\sigma}\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
@@ 5. Numerical Illustrations
In this section, we provide numerical illustrations of our results. In order to illustrate the singular control mechanism and how $(D^{*}_{t})_{t}$ ensures the nonnegativity of the wealth process, in Figure 1, we simulate the optimal primal state process $(X_{t}^{\pi^{*},c^{*}})_{t}$, the optimal dual state process $(Z^{\hat{z},D^{*}}_{t})_{t}$, and the optimal singular control $(D^{*}_{t})_{t}$. For this simulation, we fix the parameters: $r=0.03$, $\delta=0.04$, $\ell=0.6$, $\gamma=1.5$, $\kappa=0.25$, $\bar{\beta}=0.05$, $\sigma_{\beta}=0.03$, $\sigma=0.18$, and $\rho=0.6$.
@@
For this comparison, we fix the parameters: $r=0.03$, $\delta=0.04$, $\ell=0.6$, $\gamma=1.5$, $\bar{\beta}=0.05$, and $\sigma=0.18$. In the stochastic case, we additionally set $\kappa=0.25$ and $\sigma_{\beta}=0.03$, while in the constant case we set $\kappa=0=\sigma_{\beta}$ and $\beta_{0}=\bar{\beta}$ to ensure that $(\beta_{t})_{t}\equiv\bar{\beta}$. We perform the comparison for three distinct correlation parameters: $\rho\in\{-0.6,0,0.6\}$. To isolate the impact of the stochastic factor $(\beta_{t})_{t}$, we fix the same path for the Brownian motion $(W_{t})_{t}$ driving the risky asset price in both cases. For the agent who assumes a stochastic $(\beta_{t})_{t}$, we simulate $10,000$ paths of the Brownian motion $(W^{\beta}_{t})_{t}$ driving the stochastic factor, and plot the optimal wealth as the average over these paths (blue line). The red line depicts the wealth of the agent who assumes a constant $(\beta_{t})_{t}$. In the following, we denote by Stochastic Agent the agent with stochastic $(\beta_{t})_{t}$, and by Constant Agent the agent with constant $(\beta_{t})_{t}$.
@@
In all scenarios, the Stochastic Agent systematically accumulates greater long-term wealth. This outcome is driven by a universal mechanism: The Stochastic Agent’s strategic asset allocation, namely increasing exposure to the risky asset when expected returns are high (that is, when $\beta_{t}$ is high), generates substantial excess returns. As a result, the Stochastic Agent’s wealth exceeds that of the Constant Agent in the long-run. In contrast, the Constant Agent lacks this flexibility and is confined to a rigid strategy that systematically fails to exploit time-varying risk premia, leading to inferior wealth accumulation.
@@
(a) Case $\rho=-0.6$: Comparison of the optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
(b) Case $\rho=0$: Comparison of the optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
(c) Case $\rho=0.6$: Comparison of the optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
Figure 2. Comparison of optimal wealth trajectories between the Stochastic Agent and the Constant Agent ($\beta_{t}\equiv\bar{\beta}$) for different values of $\rho$. The Brownian path of $(W_{t})_{t}$ is fixed across both cases, while the Stochastic Agent’s wealth is averaged over 10,000 Brownian paths of $(W^{\beta}_{t})_{t}$.
While this long-run dominance is consistent across all regimes, short-run dynamics differ. For nonnegative correlations ($\rho=0$ and $\rho=0.6$), the advantage materializes immediately, and the Stochastic Agent maintains wealth consistently above that of the Constant Agent throughout the simulation. In contrast, the regime with $\rho=-0.6$ exhibits a distinct crossing pattern. In this case, the negative correlation provides a natural hedge, as market downturns signal improved future investment opportunities. This added safety reduces the agent’s precautionary savings motive and induces a positive income effect, effectively making the agent feel wealthier and optimally consume more in the present. The resulting higher initial consumption causes the Stochastic Agent to temporarily lag behind the Constant Agent. However, as the benefits of flexible investment decisions accumulate, the wealth paths eventually cross, confirming that the long-run advantage of adapting to $(\beta_{t})_{t}$ ultimately dominates.
@@
In Figure 3, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of labor income $\ell\in\{0.2,0.6,1\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03,\;\sigma=0.18$, and $\rho=0.6$.
@@
In Figure 4, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of risk aversion $\gamma\in\{1.2,1.5,2\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03,\;\sigma=0.18$, and $\rho=0.6$.
@@
Conversely, the consumption policy $(c^{*}_{t})_{t}$ roughly shifts upward as $\gamma$ increases. This behavior is driven by the Intertemporal Elasticity of Substitution (IES). Agents with lower risk aversion (higher IES) are more willing to postpone current consumption to capitalize on investment opportunities, whereas highly risk-averse agents (lower IES) invest less and choose to consume larger amounts of their current wealth.
@@
In Figure 5, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of the correlation parameter $\rho\in\{-0.6,0,0.6\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03$, and $\sigma=0.18$.
We observe that the consumption policy $(c_{t}^{*})_{t}$ is increasing in $\rho$. In particular, the agent with positive correlation ($\rho=0.6$) consumes the most, while the agent with negative correlation ($\rho=-0.6$) consumes the least. This pattern reflects the precautionary savings motive. When $\rho<0$, the risky asset acts as an intertemporal hedge, encouraging the agent to take a highly leveraged position. This leverage increases portfolio volatility, and to guard against risk and avoid breaching the borrowing constraint, the agent reduces current consumption. By contrast, when $\rho>0$, the asset offers weaker hedging benefits, leading to lower risky exposure, reduced volatility, and a weaker precautionary savings motive, allowing the agent to consume more.
The investment policy $(\pi_{t}^{*})_{t}$ is non-monotonic with respect to $\rho$. It rises with $\rho$ up to a wealth level of roughly $X_{t}\approx 6.5$, and then declines for higher wealth. In the high-wealth region, the policy follows standard Merton intuition: the negatively correlated agent ($\rho=-0.6$) invests more aggressively to exploit the asset’s hedging properties, while investment declines with $\rho$. At lower wealth levels ($X_{t}<6.5$), however, the pattern flips. The hedging agent’s highly leveraged strategy becomes too risky because of the threat of immediate bankruptcy, forcing sharp deleveraging. Meanwhile, the positively correlated agent ($\rho=0.6$), who naturally maintains a more conservative position, is less constrained and thus holds a higher allocation in the low-wealth region.
(a) Optimal investment policy $\pi^{*}_{t}$ for different values of correlation $\rho$.
(b) Optimal consumption policy $c^{*}_{t}$ for different values of correlation $\rho$.
Figure 5. Optimal policies for different values of correlation $\rho$.
Continuing with our cross-sectional approach, Figure 6 compares the optimal control policies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of expected excess returns $\beta_{t}\in\{0.02,0.05,0.12\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03,\;\sigma=0.18$, and $\rho=0.6$.
@@
The consumption policy $(c_{t}^{*})_{t}$ displays a non-monotonic relationship with $\beta_{t}$. At low wealth levels ($X_{t}<3$), consumption decreases as $\beta_{t}$ increases. In this region, the agent seeks to accumulate wealth more rapidly. Since a higher $\beta_{t}$ implies more profitable investment opportunities, the agent optimally cuts current consumption in order to finance larger risky positions. At higher wealth levels, however, the agent has sufficient capital to fully exploit the high expected returns. This relaxes the need for aggressive saving and allows consumption to rise, eventually exceeding the level observed in the low-$\beta_{t}$ state.
@@ 6. Conclusions
In this paper, we have studied the infinite-horizon consumption and portfolio problem of an investor subject to a no-borrowing constraint and labor income, in a market where expected excess returns follow a mean-reverting Ornstein-Uhlenbeck process. Using a Lagrange duality approach, we have formulated the dual problem as a two-dimensional singular control problem involving the marginal value of wealth and the stochastic factor. The solution is governed by an auxiliary optimal stopping problem, which features a free boundary separating the continuation and stopping regions. We have provided a detailed probabilistic analysis for this optimal stopping problem and established properties of the free boundary and of the value function. Finally, we have retrieved the solutions to the primal optimization problem via duality and provided numerical illustrations.

### 11 Mar 2026 v3

@@ 4. Back to the Primal Problem
The process $\tau^{D}:=\{\tau^{D}(\alpha)\mid\alpha\in(0,1]\}$ has nonincreasing, left-continuous sample paths and hence it admits right-limits $\tau^{D}_{+}(\alpha):=\inf\{t\geq 0\mid D_{t}\leq\alpha\}$. The set of points $\alpha\in(0,1]$ at which $\tau^{D}(\alpha)(\omega)\neq\tau^{D}_{+}(\alpha)(\omega)$ is countable for a.e. $\omega\in\Omega$. Since $(D_{t})_{t}$ is right-continuous and $\tau^{D}(\alpha)$ is the first entry time of an open set, it is an $\mathcal{F}^{W}_{t+}$-stopping time for any given and fixed $\alpha\in(0,1]$. However, $(\mathcal{F}^{W}_{t})_{t}$ is right-continuous, hence $\tau^{D}(\alpha)$ is an $\mathbb{F}^{W}$-stopping time. For a fixed $z>0$ and $y\in(0,z]$, we consider $\alpha=\frac{y}{z}$ in the following. Upon using sub-optimality of $\tau^{D}(\frac{y}{z})$ for (3.13), we have
@@
where we used that $D^{*}\in\mathcal{D}$ since $(\xi_{t}^{*})_{t}$ is nondecreasing, càdlàg, and $\mathbb{F}^{W}$-adapted.
@@
$c_{t}^{*}=(Z_{t}^{\hat{z},\beta,D^{*}})^{-\frac{1}{\gamma}}\qquad\text{and}\qquad\pi_{t}^{*}=\frac{\beta_{t}}{\sigma^{2}}Z_{t}^{\hat{z},\beta,D^{*}}\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})+\frac{\sigma_{\beta}}{\sigma}\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
@@
$=\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\big(u(c_{t}^{*})-\tilde{u}(Z_{t}^{D^{*}})\big)\,dt\bigg]-\mathbb{E}_{\hat{z},\beta}\bigg[\int_{0}^{\infty}e^{-\delta t}\ell Z_{t}^{D^{*}}\,dt\bigg].$
By Proposition 2.4-(2), there exists an investment strategy $\pi^{*}$ such that $(\pi^{*},c^{*})\in\mathcal{A}(x)$ thanks to $\chi(\hat{z})=x$. Hence, by using (4.14), we obtain
@@
so that $c_{t}^{*}=(Z_{t}^{\hat{z},\beta,D^{*}})^{-\frac{1}{\gamma}}$ is optimal. Moreover, by the strong Markov property and $x=-\tilde{V}_{z}(\hat{z},\beta)$, we have
@@
$-\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\left[\kappa(\overline{\beta}-\beta_{t})\,dt-\sigma_{\beta}\,dW_{t}\right]$
@@
$-\frac{1}{2}\sigma_{\beta}^{2}\tilde{V}_{z\beta\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\,dt-\frac{\beta_{t}}{\sigma}Z_{t}^{\hat{z},\beta,D^{*}}\sigma_{\beta}\tilde{V}_{zz\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})\,dt,\quad t\geq 0.$
Comparing the $dW_{t}$ terms in (2.3) and (4.19) finally gives
$\pi_{t}^{*}=\frac{\beta_{t}}{\sigma^{2}}Z_{t}^{\hat{z},\beta,D^{*}}\tilde{V}_{zz}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t})+\frac{\sigma_{\beta}}{\sigma}\tilde{V}_{z\beta}(Z_{t}^{\hat{z},\beta,D^{*}},\beta_{t}),\quad t\geq 0,$
@@ 5. Numerical Illustrations
In this section, we provide numerical illustrations of our results. In order to illustrate the singular control mechanism and how $(D^{*}_{t})_{t}$ ensures the nonnegativity of the wealth process, in Figure 1, we simulate the optimal primal state process $(X_{t}^{\pi^{*},c^{*}})_{t}$, the optimal dual state process $(Z^{\hat{z},D^{*}}_{t})_{t}$, and the optimal singular control $(D^{*}_{t})_{t}$. For this simulation, we fix the parameters: $r=0.03$, $\delta=0.04$, $\ell=0.6$, $\gamma=1.5$, $\kappa=0.25$, $\bar{\beta}=0.05$, $\sigma_{\beta}=0.03$, and $\sigma=0.18$.
@@
For this comparison, we fix the parameters: $r=0.03$, $\delta=0.04$, $\ell=0.6$, $\gamma=1.5$, $\bar{\beta}=0.05$, and $\sigma=0.18$. In the stochastic case, we additionally set $\kappa=0.25$ and $\sigma_{\beta}=0.03$, while in the constant case we set $\kappa=0=\sigma_{\beta}$ and $\beta_{0}=\bar{\beta}$ to ensure that $(\beta_{t})_{t}\equiv\bar{\beta}$. In both cases, we simulate $10,000$ paths of the Brownian motion $(W_{t})_{t}$, and plot the optimal wealth as the average over these paths. The blue line depicts the wealth of the agent who assumes a stochastic $(\beta_{t})_{t}$, while the red line represents the wealth of the agent who assumes a constant $(\beta_{t})_{t}$. In the following, we denote by Stochastic Agent the agent with stochastic $(\beta_{t})_{t}$, and by Constant Agent the agent with constant $(\beta_{t})_{t}$.
@@
We observe that the Stochastic Agent systematically accumulates greater wealth. This outcome is driven by a robust economic mechanism: The Stochastic Agent’s strategic asset allocation, namely increasing exposure to the risky asset when expected returns are high (that is, when $\beta_{t}$ is high), generates substantial excess returns. As a result, the Stochastic Agent’s wealth exceeds that of the Constant Agent. In contrast, the Constant Agent lacks this flexibility and is confined to a rigid strategy that systematically fails to exploit time-varying risk premia, leading to inferior wealth accumulation.
@@
Figure 2. Comparison of the average optimal wealth trajectories for the Stochastic Agent and the Constant Agent ($(\beta_{t})_{t}\equiv\bar{\beta}$).
At the same time, the initial advantage builds up only gradually, so that the two wealth paths remain relatively close at the beginning of the horizon. Economically, under perfect negative correlation, adverse contemporaneous shocks to the risky asset are accompanied by improvements in future investment opportunities. This hedging effect dampens the short-run urgency of aggressive portfolio adjustment and smooths the initial wealth dynamics. Hence, although the Stochastic Agent is still able to exploit the time-varying investment opportunity set and eventually achieves substantially higher wealth, the momentum of this outperformance is weaker at the start. Over time, however, the cumulative benefits of adapting consumption and portfolio decisions to the stochastic factor $(\beta_{t})_{t}$ become increasingly important, leading to a persistent and widening wealth advantage over the Constant Agent.
@@
In Figure 3, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of labor income $\ell\in\{0.2,0.6,1\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03$, and $\sigma=0.18$.
@@
In Figure 4, we compare the optimal control strategies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of risk aversion $\gamma\in\{1.2,1.5,2\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03$, and $\sigma=0.18$.
@@
Conversely, the consumption policy $(c^{*}_{t})_{t}$ shifts upward as $\gamma$ increases. This behavior is driven by the Intertemporal Elasticity of Substitution (IES). Agents with lower risk aversion (higher IES) are more willing to postpone current consumption to capitalize on investment opportunities, whereas highly risk-averse agents (lower IES) invest less and choose to consume larger amounts of their current wealth.
@@
Continuing with our cross-sectional approach, Figure 5 compares the optimal control policies $(\pi^{*}_{t})_{t}$ and $(c^{*}_{t})_{t}$ for different values of expected excess returns $\beta_{t}\in\{0.02,0.05,0.12\}$, while we fix the other parameters as: $r=0.03,\;\delta=0.04,\;\ell=0.6,\;\gamma=1.5,\;\kappa=0.25,\;\bar{\beta}=0.05,\;\sigma_{\beta}=0.03$, and $\sigma=0.18$.
@@
The consumption policy $(c_{t}^{*})_{t}$ displays a non-monotonic relationship with $\beta_{t}$. At low wealth levels ($X_{t}<3.5$), consumption decreases as $\beta_{t}$ increases. In this region, the agent seeks to accumulate wealth more rapidly. Since a higher $\beta_{t}$ implies more profitable investment opportunities, the agent optimally cuts current consumption in order to finance larger risky positions. At higher wealth levels, however, the agent has sufficient capital to fully exploit the high expected returns. This relaxes the need for aggressive saving and allows consumption to rise, eventually exceeding the level observed in the low-$\beta_{t}$ state.
@@ 6. Conclusions
In this paper, we have studied the infinite-horizon consumption and portfolio problem of an investor subject to a no-borrowing constraint and labor income in the complete-market case of the Kim–Omberg model, where expected excess returns follow a mean-reverting Ornstein–Uhlenbeck process and are perfectly negatively correlated with the risky asset return shocks. Using a Lagrange duality approach, we have formulated the dual problem as a two-dimensional singular control problem involving the marginal value of wealth and the stochastic factor. The solution is governed by an auxiliary optimal stopping problem, which features a free boundary separating the continuation and stopping regions. We have provided a detailed probabilistic analysis for this optimal stopping problem and established properties of the free boundary and of the value function. Finally, we have retrieved the solutions to the primal optimization problem via duality and provided numerical illustrations.
An interesting direction is the extension to the incomplete-market case. In that setting, the duality approach developed here is no longer directly applicable, as our proof of strong duality would break down, and a different mathematical approach, likely involving the associated HJB equation and verification arguments, would be needed. Such a challenging extension is left for future research.